OpenStudy (anonymous):

Use the commutative and associative properties as needed to simplify the expression: 7/8(8/7x)

OpenStudy (anonymous):

Please help and explain

OpenStudy (zzr0ck3r):

\[\dfrac{7}{8}(\dfrac{8}{7}x)\]?

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

And thanks for helping

OpenStudy (zzr0ck3r):

\[\dfrac{7}{8}(\dfrac{8}{7}x)=\dfrac{7}{8}(\dfrac{8}{7})x=\dfrac{\cancel{7}}{\cancel{8}}(\dfrac{\cancel{8}}{\cancel{7}})x=x\]

OpenStudy (anonymous):

So x just equals x?

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

@zzr0ck3r nice use of \(\cancel{\xi}\)!!

OpenStudy (anonymous):

I was confused, but the answer is X=X, correct?

OpenStudy (zzr0ck3r):

What you started with equals \(x\).

OpenStudy (zzr0ck3r):

I love the latex code that is obvious:)

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

Thank you